Hsh13:

You take out a cash advance of $2,790.00 on a credit card with a 3.5% cash advance fee and a 13.5% APR. What is the amount due by the end of the month? $3,277.48 $2,920.14 $2,887.65 $2,821.39

Vocaloid:

@sillybilly123 if it wouldn't be too much trouble would you mind taking a look at this?

sillybilly123:

Three components to the end of month balance - the principal advanced = 2,790 - the cash advance fee = 3.5% * 2,790 - the APR (cost of borrowing). For APR, you need to convert the 13.5% APR into a monthly rate \(r\) For a nominal APR, more common in the US, it's just \(r = \dfrac{13.5}{12} = 1.125 \%\) More sophisticated models us an effective APR which needs to be de-compounded: \((1 + r)^{12} = 1 + APR \implies r = (1 + APR)^{1/12} - 1 = 1.06 \%\) Using the nominal APR, the cost due to APR is: 2,790 * 1.125 % Adding it all up: 2,790 + 3.5 % * 2,790 + 2,790 * 1.125 % = 2,790 ( 1 + 3.5 % + 1.125 %) = 2,919.04